# Scalable Private Search with Wally

### v5

@@ Abstract
This paper presents Wally, a private search system that supports efficient semantic and keyword search queries against large databases. When sufficiently many clients are making queries, Wally’s performance is significantly better than previous systems. In previous private search systems, for each client query, the server must perform at least one expensive cryptographic operation per database entry. As a result, performance degraded proportionally with the number of entries in the database.
In Wally, we remove this limitation. Specifically, for each query the server performs cryptographic operations against only a few database entries. We achieve these results by requiring each client to add a few fake queries, and send each query via an anonymous network to the server at independently chosen random instants. Additionally, each client also uses somewhat homomorphic encryption (SHE) to hide whether a query is real or fake. Wally provides $(\epsilon,\delta)$-differential privacy guarantee, which is an accepted standard for strong privacy.
@@
The number of fake queries each client makes depends inversely on the number of clients making queries. Therefore, the fake queries’ overhead vanishes as the number of clients increases, enabling scalability to millions of queries and large databases. Concretely, Wally can process eight million queries in 117 minutes, or just under two hours. That is around four orders of magnitude faster than the state of the art.
1 Introduction
@@
Consider a scenario where a client holding a search query wants to retrieve relevant information from a large database hosted on a server. This scenario reflects the flow of current search engines such as Google, Bing, and DuckDuckGo. These search engines improve the quality of our lives by promptly providing helpful information. Nevertheless, they require learning the client’s query in order to respond accurately, hence providing no query privacy. In some situations, learning the query could reveal sensitive and personal information about the client. For example, consider the queries “High blood pressure with late-stage HIV?”, “Closest diabetes clinic from Albuquerque”, and “Best divorce lawyers”. Revealing these queries to search engines could be damaging to the client. However, even if the queries do not appear sensitive, they may allow the server to learn personal information like medical information, marital status, sexuality, etc. Therefore, enabling query privacy is critical for search engines.
@@
Private search engines opt to provide query privacy. However, these systems have significantly higher computational and communication overhead than the state-of-the-art insecure solutions. For example, in Tiptoe [25], a state-of-the-art private search engine, a single client query is 21 MB and requires 339 core-seconds of server computing against a database of $400$ million entries. The high computational cost is because, for each client query, the server must scan the entire database; otherwise, it will learn which database entries the client is not interested in. Similarly, the client must send cryptographic material for each database entry, which results in high communication cost [^1] In reality, these systems trade off response size with request size by using a standard database clustering trick. We detail this trade off in a later section.. An additional limitation of Tiptoe is that it requires each client to store a database-dependent state of around 50 MB for each client query.
In this paper, we present Wally, a privacy preserving search engine that efficiently scales to the performance constraints of large-scale search engines. Wally supports semantic search, where the client wants to fetch the most relevant documents to a query, and keyword search, where the client wants to retrieve documents associated with a private keyword. To provide privacy, our key observation is that in search engines, there are always enough clients making queries. We exploit this observation to gain privacy by devising a way to hide a particular query in the crowd of queries. In our protocol, the server computation and the communication are amortized over the number of clients making the requests; therefore, when the number of clients is large, the per-query cost becomes minimal.
Wally achieves these results by relaxing its privacy notion to $(\epsilon,\delta)$-differential privacy. In Wally, the server learns a noised differentially private histogram of anonymized queries from all clients, over a partition of the server database. We emphasize that previous works, e.g., Tiptoe, are fully oblivious. That is, the server learns nothing besides that a particular client is making a query. Therefore, the privacy guarantee of schemes like Tiptoe is stronger than Wally. However, differential privacy is an accepted standard for strong privacy [15]. Specifically, every client in Wally is guaranteed that any statistical inferences are insensitive to their data. In other words, their participation does not significantly alter the server’s view, except with probability $\delta$, and thus provides strong privacy protection.
At a high level, Wally works in epochs. Within each epoch, a fixed number of users make queries. The epoch length is picked to ensure enough clients are present to make queries, and their queries will be processed by the end of an epoch. The server divides the database into $K$ clusters. Each query consists of a relevant cluster index, and the server response is all the entries in the cluster. The queries arrive at the server via an anonymous network, ensuring all identifiable information like IP address is removed from the queries. The server then generates a response to each query independently and sends it back to the respective client via an anonymous network.
@@
Even though queries are anonymized, they do not provide query privacy. The server can infer information about the client query by exploiting the access pattern (if a particular cluster is accessed) or arrival time (when the server receives queries). To prevent this leakage, Wally implements the following two changes at the client.
First, to ensure that the server does not learn information by observing the access patterns, each client in Wally makes fake queries. For each fake query, the client picks a randomly selected bucket. This ensures that the server cannot differentiate between real and fake queries. The number of fake queries is sampled from a particular distribution to ensure the revealed histogram of queries is $(\epsilon,\delta)$-differentially private. In expectation, each client makes $O(\frac{C\Delta\log(1/\delta)}{U\epsilon})$ fake queries where $U$ is the number of clients within an epoch, and $\Delta$ is the maximum number of queries each client can make in an epoch. Note that increasing the number of clients decreases the average number of fake queries. The overhead due to fake queries can be traded with privacy because more fake queries are required as $\epsilon$ decreases. We emphasize that fake queries do not affect the accuracy or correctness seen by any client, as the client can discard the responses due to these queries.
@@
Second, Wally also ensures that arrival time and order of the queries leak no information about the client to the server. To achieve this, instead of making a query instantly, the clients make queries at random one second slots within an epoch. Also, an anonymous network collects the requests received within a particular slot and sends them as a batch to the server.
Anonymizing the queries, adding fake queries that are indistinguishable from real ones, and randomizing query order and time guarantees differential privacy.
@@
As mentioned above, for each query, the server response contains all the entries in the cluster. For large clusters, this would result in very high response overhead. In Wally, we utilize lattice-based, somewhat homomorphic encryption (SHE) to reduce the response overhead. Specifically, for semantic search, each query additionally includes a semantic information (dependent on the client’s input) encrypted under SHE. The server computes the distance function between the encrypted information and the cluster entries under SHE and returns encrypted scores. This reduces the response because the size of encrypted scores is significantly smaller than the entries.
For a keyword search, the client sends an encrypted PIR query, and the server runs an SHE-based keyword private information retrieval (PIR) protocol within the cluster, which returns data associated with the client keyword. To ensure that the server cannot differentiate between real and fake queries, for every fake query the client includes encrypted semantic, or keyword queries based on fake data. We propose various optimizations that significantly improve performance for both kinds of searches.
Overall, for a database with 35 million entries Wally can process eight million real queries in 117 minutes on 320 cores, a drastic improvement over the state-of-the-art, which would have required years to process the same queries. Additionally, the size of each request is 53 times smaller and the size of each response is 3.5 times smaller compared to Tiptoe.
Our contributions are as follows:
  • We describe Wally’s design, which enables privacy-preserving large-scale search engines. Its novelty stems from the combination of anonymity, fake queries added by the clients, and the client’s use of SHE to encrypt its queries.
  • We also describe the implementation of Wally, showing efficiency and scalability along with a robust privacy guarantee. Generally, SHE computations have a high performance overhead. To achieve desired performance constraints, we propose several optimizations to SHE computation (distance computation and PIR) that could be of independent interest.
  • We evaluate Wally for private image search, demonstrating both strong privacy and low overhead.
Limitation.
Wally requires a set number of clients always present in an epoch to make queries. This assumption aligns with current search engines. For example, hundreds of millions of people make queries in Google search daily [18]. We remark that for systems with few users, the overhead due to fake queries could get quite high, therefore fully oblivious protocols, like Tiptoe, for these systems would be more suitable.
2 Background
@@
Let $[n]:=\{1,\dots,n\}$ be the set of the first $n$ natural numbers starting from one. Matrices and vectors are represented as capital boldface letters and capital lowercase, respectively, e.g., $\mathbf{M}$ and $\mathbf{v}$. The entry-wise, Hadamard, product of two vectors is denoted as $\mathbf{v}\odot\mathbf{v}^{\prime}$. All logarithms are the natural log unless noted otherwise. We denote a random variable being sampled from a probability distribution $D$ as $a\leftarrow D$. All distributions in this work are over finite sets and $a\leftarrow S$ denotes uniformly sampling from $S$ when $S$ is a discrete set. We denote the integers as $\mathbb{Z}$ and the integers modulo a positive integer $Q$ as $\mathbb{Z}_{Q}$. Let $\mathcal{NB}(r,p)$ denote the negative binomial distribution parameterized by $r,p$. Its probability mass function is $Pr(X=k)=\binom{k+r-1}{k}(1-p)^{k}p^{r}$, $X\sim\mathcal{NB}(r,p)$, for every non-negative integer $k$. An important fact about the negative binomial distribution is that it is infinitely divisible, Definition 1.
@@
A distribution $\mathcal{D}$ is infinitely-divisible if for all $n\in\mathbb{N}$, it can be expressed as the sum of $n$ i.i.d. variables. That is, there exists a distribution $\mathcal{D}_{n}$ such that $\mathcal{X}_{1}+\cdots+\mathcal{X}_{n}\sim\mathcal{D}$ and $\mathcal{X}_{i}\sim\mathcal{D}_{n}$.
2.1 (Somewhat) Homomorphic Encryption
@@
In Table 1 we show concrete run times and noise growth for BFV, an efficient SHE scheme. In general, addition and plaintext multiplication are efficient and addition has minimal noise growth. The latter has larger growth proportional to $t$. Ciphertext-ciphertext multiplication has large noise growth and is somewhat inefficient since it requires multiplying the ciphertext polynomials over the integers then rounding down back to $\mod Q$. Ciphertext rotation adds little noise but it is somewhat slow due to key-switching. Therefore, minimizing the number of ciphertext multiplications and ciphertext rotations is crucial for scalable applications.
@@
Ciphertext rotation and multiplication require evaluation keys that depend on a secret key. Additionally, each rotation requires a separate evaluation key. Usually, these keys can be reused, but as mentioned later, in Wally  the client must send fresh keys for each query. Therefore, these keys contribute to the client’s request size.
@@
Table 1: Experimental computation cost and noise growth of each BFV homomorphic operation. The polynomial degree $n$ is $4096$, the ciphertext modulus $Q$ has $83$ bits, and the plaintext modulus $t=40961$ (16 bits). $\mathsf{CtRotate}$ noise is measured on repeated rotations and the others are from two input ciphertexts with the same noise levels. Time costs are measured with the swift-homomorphic-encryption library [6] (git commit b70d927) on an Intel Xeon w3-2423 using a single core.
@@
We use the standard notion of differential privacy (DP) [14, 13]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private if for all $\epsilon\geq 0$ and $\delta\in[0,1]$, any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$. Let $\zeta$ be an output of a mechanism $M$, then the privacy loss in observing $\zeta$ between two datasets $X,X^{\prime}$ as input is
@@
Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations. We use basic composition in DP [35, Lemma 2.3].
@@
Anonymous networks (ANs), also called anonymization networks and anonymous communication protocols, are protocols which offer user anonymity by using cryptographic techniques such as onion routing [22], or mixnets [11]. Throughout the paper, we assume that ANs 1) strip all identifying information from incoming messages from clients, and 2) permute all received messages within a batch of messages from clients before sending them to the server.
@@
Though the paper is mostly presented for semantic search, Wally applies to keyword search as well by using keyword variant of private information retrieval (PIR). PIR protocols are cryptographic client-server protocols where a server holds a large database and the client can privately retrieves an entry corresponding to a keyword from the database. In this paper, we focus on protocols based on a single server [29]. There is a also notion of differentially-private PIR in the multi-server setting [33, 2], where the database is stored across non-colluding servers while guaranteeing differential privacy. We can not directly use these protocols in Wally which stores the database on one server or multiple servers controlled by a single entity.
3 System Overview

### v6

@@ Abstract
This paper presents Wally, a private search system that supports efficient search queries against large databases. When sufficiently many clients are making queries, Wally’s performance is significantly better than previous systems while providing a standard privacy guarantee of $(\epsilon,\delta)$-differential privacy.
Specifically, for a database with 3.2 million entries, Wally’s queries per second (QPS) is 7-28x higher, and communication is 6.69-31x smaller than Tiptoe, a state-of-the-art private search system. In Wally, each client adds a few fake queries and sends each query via an anonymous network to the server at independently chosen random instants. We also use somewhat homomorphic encryption (SHE) to reduce the communication size.
@@
The number of fake queries each client makes depends inversely on the number of clients making queries. Therefore, the overhead of fake queries vanishes as the number of honest clients increases, enabling scalability to millions of queries and large databases.
I Introduction
@@
Consider a scenario where a client holding a potentially private search query wants to search information from a large database hosted on a server. This scenario reflects the flow of current search engines such as Google, Bing, and DuckDuckGo. These search engines improve the quality of our lives by promptly providing helpful information. Nevertheless, they require learning the client’s query in order to respond accurately, hence providing no query privacy. In some situations, learning the query could reveal sensitive and personal information about the client. For example, consider the queries “High blood pressure with late-stage HIV?”, “Closest diabetes clinic from Albuquerque”, and “Best divorce lawyers”. Revealing these queries to search engines could be damaging to the client. However, even if the queries do not appear sensitive, they may allow the server to learn personal information like medical information, marital status, sexuality, etc. Therefore, enabling query privacy is critical for search engines.
Private search systems prioritize user privacy. There is a long list of works that have explored protecting query privacy, by periodically injecting fake queries that resemble real queries [1, 2, 3, 4, 5] or obfuscating users’ original queries using decoy terms [6, 7, 8, 9]. These approaches are quite efficient. However, they lack a standard privacy guarantee and often rely on assumptions that limit adversarial capabilities. Many systems assume adversaries have restricted knowledge—whether unaware of the obfuscation algorithm, lacking background data, or using a predefined inference strategy. As a result, when these assumptions are removed or a more capable adversary is considered, fake queries become identifiable, ultimately breaking query privacy [10, 11, 12, 13].
@@
Recent privacy search systems do provide the strongest privacy guarantee - the server does not learn any information about the client query irrespective of knowledge and capability, but these schemes incur very high server computation and communication which makes them impractical. For example, Tiptoe [14], a state-of-the-art private search system, requires a communication of around 17 MB and provides queries per second (QPS) of only 909, with 10,000 cores, for a database with 3.2 million entries. The high computational cost is because, for each client query, the server must scan the entire database; otherwise, it will learn which database entries the client is not interested in. Similarly, the client must send cryptographic material for each database entry, which results in a high communication cost [^1] In reality, these systems trade off response size with request size by using a standard database clustering trick. We detail this trade-off in a later section..
The world of private search thus is in an undesirable state. On one hand, traditional approaches do not provide a standard privacy guarantee, making them vulnerable; on the other hand, schemes that provide a strong privacy guarantee have very high computational and communication requirements, due to which they cannot be used in large-scale systems.
Our system.
We present Wally, a privacy preserving search system that provides standard privacy guarantees while maintaining efficiency. Our key observation is that in search engines, there are always enough clients making queries. Using this observation we explore a new trade-off between efficiency and privacy. Specifically, in Wally some small information is leaked to the server, but this leakage is proven to preserve differential privacy for the participating users. In return, Wally’s computational and communication overhead reduces significantly compared to the state of the art. Specifically, for a database with 3.2 million entries, Wally’s QPS is 7$\sim$28x higher and communication is 6.69$\sim$31x smaller than Tiptoe. Additionally, the search quality (presented as MRR@@@@100 scores) of Wally is decently better than Tiptoe.
Wally relaxes its privacy notion to $(\epsilon,\delta)$-differential privacy. The privacy guarantee of schemes like Tiptoe is stronger than Wally. However, differential privacy is an accepted standard for strong privacy [15]. Intuitively, every client in Wally is guaranteed that leakage to the server does not change by too much when their input is modified or removed. More formally, the server’s ability to infer their queries does not improve by more than a factor of $e^{\epsilon}$ , except with probability at most $\delta$.
We emphasize that we do not make any other assumption about the client queries or the adversary’s capabilities. Additionally, the privacy guarantee only gracefully degrades when the adversary sees query history.
Summary of techniques.
Like Tiptoe, Wally also uses standard search techniques based on semantic embeddings. At a high level, the embedding model maps any text string into a floating-point vector of dimension $d$ with a semantic guarantee: embeddings corresponding to similar texts will have a high inner product score (cosine similarity). This way, the search problem is reduced to the problem of private nearest neighbors: The client must find the embedding vector that maximizes the inner product with the query vector.
The server divides the database into $K$ clusters and sends cluster centroids to each client. The system works in epochs. The epoch length is picked to ensure that enough honest clients are present to make queries and there are enough server resources to process all the queries by the end of an epoch. Within each epoch, a large number of honest users make queries. We discuss this assumption in more detail towards the end of the introduction.
@@
For each query, the client finds $\Delta$ closest clusters and forwards them to the server. The server responds by sending all the embeddings in those clusters. The queries arrive at the server via an anonymous network, ensuring that all identifiable information, such as IP address, is removed from them. The server then generates a response to each query independently and sends it back to the respective client via an anonymous network. Although queries are anonymized, this does not guarantee query privacy. The server can infer information about the client’s query by exploiting the access pattern (if a particular cluster is accessed) or arrival time (when the server receives queries). To prevent this, Wally implements the following two changes at the client.
First, each client in Wally makes fake queries. To generate a single fake query, the client picks a random cluster. The client does not have to interact with the server to generate random queries and generates them independently. The number of fake queries is sampled from a negative binomial distribution to ensure $(\epsilon,\delta)$-differential privacy. Looking ahead, to maintain privacy, the client must send only a single cluster per query. We emphasize that fake queries do not affect the accuracy or correctness seen by any client, as the client can discard the responses due to these queries.
@@
Second, Wally also ensures that neither the arrival time nor the order of the queries leak any information about the client to the server. To achieve this, instead of sending a query instantly, each client sends queries in randomly sampled one-second slots within an epoch. Additionally, an anonymous network collects the requests received within a particular slot and sends them as a batch to the server.
We show that anonymizing the queries, adding fake queries, and randomizing the query schedule is equivalent to sharing a noised histogram of queries from all the clients with the server. We then prove that this histogram guarantees $(\epsilon,\delta)$ privacy.
@@
As mentioned above, for each query, the server response contains all the entries in the cluster. For large clusters, this would result in a high response overhead. We utilize lattice-based, homomorphic encryption (SHE) to reduce the response overhead. Specifically, for each real query, the client freshly encrypts the query embedding under SHE and sends it along with the corresponding cluster to the server. For a fake query, the client encrypts just a zero embedding. To compute the distance score, the server computes the dot-product between the encrypted embeddings and the cluster embeddings under SHE. This reduces the response because the size of encrypted distance scores is significantly smaller than the embedding size.
Formal contributions. Our contributions are as follows:
  • An efficient private search system, named Wally. A novel system that meticulously achieves $(\epsilon,\delta)$ differential privacy for client queries. It achieves this by carefully stitching various steps to enhance privacy. Moreover, we provides a formal security proof that demonstrates the protocol’s security against a semi-honest adversary.
  • Novel optimizations enable efficient computation of distance scores under SHE.
  • A standalone open-source Swift library that includes an implementation of the BFV SHE scheme and various privacy primitives, including Private dot-product and Keyword Private Information Retrieval. This library is highly versatile and can be easily integrated into various privacy applications.
  • We compare Wally with two state-of-art schemes and demonstrate how Wally provides a better trade-off between privacy and performance.
Epoch size in Wally.
Wally works in epochs, and we assume there are enough clients present in an epoch. Note that large epochs allow considering more clients. This assumption aligns with current search engines. For example, hundreds of millions of people make queries in current search systems daily [16]. So we could consider a day-size epoch with millions of users. However, note that epoch size also dictates the client experience; long epochs mean some clients will have to wait till the end of epoch to get response to their queries. Picking the very short epoch is not desirable as well. As few clients will present, this means high overhead due to fake queries. Concretely, within an epoch, on average, each client makes $O(\frac{C\Delta\log(1/\delta)}{U\epsilon})$ fake queries, where $U$ is the number of clients within an epoch, and $C$ is the number of clusters on the server. Moreover, $\Delta$ is the maximum number of clusters each client can probe in an epoch.
Our experiments assume epochs of length 1-10 minutes, with that at least 100,000 to 500,000 clients are present to make queries within any epoch. Malicious clients can affect privacy by not submitting the fake queries. In our privacy calculation, we assume that at least half the clients are honest and exclude fake queries from malicious clients.
We remark that Tiptoe does not have this requirement. Therefore, for systems with few clients (like a few hundred users), the overhead from fake queries will be relatively high; therefore, for such a situation, Tiptoe would be more suitable.
We further discuss these assumptions in Section VIII.
II Background
@@
Let $[n]:=\{1,\dots,n\}$ be the set of the first $n$ natural numbers starting from one. Matrices and vectors are represented as capital boldface letters and capital lowercase, respectively, e.g., $\mathbf{M}$ and $\mathbf{v}$. The entry-wise, Hadamard, product of two vectors is denoted as $\mathbf{v}\odot\mathbf{v}^{\prime}$. All logarithms are the natural log unless noted otherwise. We denote a random variable $a$ being sampled from a probability distribution $D$ as $a\leftarrow D$. All distributions in this work are over finite sets and $a\leftarrow S$ denotes uniformly sampling from $S$ when $S$ is a discrete set. We denote the integers as $\mathbb{Z}$ and the integers modulo a positive integer $Q$ as $\mathbb{Z}_{Q}$. Let $\mathcal{NB}(r,p)$ denote the negative binomial distribution parameterized by $r,p$. Its probability mass function is $Pr(X=k)=\binom{k+r-1}{k}(1-p)^{k}p^{r}$, $X\sim\mathcal{NB}(r,p)$, for every non-negative integer $k$. An important fact about the negative binomial distribution is that it is infinitely divisible, Definition 1.
@@
A distribution $\mathcal{D}$ is infinitely divisible if for all $n\in\mathbb{N}$, it can be expressed as the sum of $n$ i.i.d. variables. That is, there exists a distribution $\mathcal{D}_{n}$ such that $\mathcal{X}_{1}+\cdots+\mathcal{X}_{n}\sim\mathcal{D}$ and $\mathcal{X}_{i}\sim\mathcal{D}_{n}$.
II-A (Somewhat) Homomorphic Encryption
@@
In Table I we show concrete run times and noise growth for BFV, an efficient SHE scheme. In general, addition and plaintext multiplication are efficient and addition has minimal noise growth. The latter has larger growth proportional to $t$. Ciphertext-ciphertext multiplication has large noise growth and is somewhat inefficient since it requires multiplying the ciphertext polynomials over the integers then rounding down back to $\mod Q$. Ciphertext rotation adds little noise but it is somewhat slow due to key-switching. Therefore, minimizing the number of ciphertext multiplications and ciphertext rotations is crucial for scalable applications.
@@
BFV multiplication, rotation, and substitution operations require evaluation keys that depend on the client’s secret key. The server could store the evaluation keys once in typical SHE applications and reuse them across various client requests. However, the server learns the requests belonging to the same client. Looking ahead, this will break the privacy guarantee in Wally  therefore, the client must send fresh evaluation keys with each request. Therefore, these keys contribute to the client’s request size.
@@
TABLE I: Experimental computation cost and noise growth of each BFV homomorphic operation. The polynomial degree $n$ is $4096$, the ciphertext modulus $Q$ has $83$ bits, and the plaintext modulus $t=40961$ (16 bits). $\mathsf{CtRotate}$ noise is measured on repeated rotations and the others are from two input ciphertexts with the same noise levels. Time costs are measured with the swift-homomorphic-encryption library [19] (git commit b70d927) on an Intel Xeon w3-2423 using a single core.
0
@@
We use the standard notion of differential privacy (DP) [20, 21]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private for $\epsilon\geq 0$ and $\delta\in[0,1]$, if for any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$. Let $\zeta$ be an output of a mechanism $M$, then the privacy loss in observing $\zeta$ between two datasets $X,X^{\prime}$ as input is
@@
Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations. We use basic composition in DP [22, Lemma 2.3].
@@
Anonymous networks (ANs), also called anonymization networks and anonymous communication protocols, are protocols which offer user anonymity by using cryptographic techniques such as onion routing [23], or mixnets [24]. Throughout the paper, we assume that ANs 1) strip all identifying information from incoming queries from clients, and 2) collect all of them received in a second and send them in a batch to the server.
@@
Though the paper is mostly presented for semantic search, Wally applies to keyword search as well by using keyword variant of private information retrieval (PIR). PIR protocols are cryptographic client-server protocols where a server holds a large database and the client can privately retrieves an entry corresponding to a keyword from the database. In this paper, we focus on protocols based on a single server [25]. There is also a notion of differentially-private PIR in the multi-server setting [26, 27], where the database is stored across non-colluding servers while guaranteeing differential privacy. We can not directly use these protocols in Wally which stores the database on one server or multiple servers controlled by a single entity.
II-B Threat Model
We break Wally’s entities into three groups: clients, the anonymous network’s (AN) hops, and the server. The server and AN hops are modeled as semi-honest adversaries attempting to learn a client’s query. Clients can behave maliciously, intentionally deviating from the protocol, in an attempt to learn an honest client’s query. The server’s database is assumed to be public.
Assumptions.
  • Clients: We assume there are at least two honest clients. Note that malicious client could break privacy of the protocol by looking ahead, as described in Theorem 1 and in VIII. Wally needs many more honest clients for better performance, as looking ahead since Wally relies on differential privacy. Malicious clients can collude with each other, an anonymous network hop, or the server. If they collude with an anonymous network hop, then we assume they collude with the server and a hop that is not the first hop.
  • Anonymous network hops: We assume there are at least two hops, and that at least one hop is honest. This is a standard assumption in anonymous networks. For example, oblivious HTTP [28] has two hops where the second hop is assumed to be operated by the server and the first hop is assumed to be operated independently by a different party.
  • Server: The server can collude with any anonymous network hop besides the entry hop, as well as clients. We assume the entry hop cannot collude with the server since the server would learn which client sent which queries, fake and real.
  • Cryptographic assumptions: Lastly, we assume the CPA security of RLWE-based SHE [29, 17].
Client privacy.
Under the above assumptions, Wally protects client query utilizing two privacy notions:
  • The client’s queries in Wally are protected under CPA-security of Wally’s SHE scheme.
  • Which clusters the client queries is protected under a differential privacy guarantee, Theorem 1.
Overall, Wally provides a differential privacy guarantee on the client’s queries because of the latter bullet above. Further, we make no assumptions on the server’s knowledge of the expected query distribution across all clients.
III Private search with inefficient communication